OpenStudy (anonymous):

PHYSICS. Help me solve this please A plane flying due north at 100 m/s is blown by a 500 m/s strong wind due east. what is the plane resultant velocity?

OpenStudy (anonymous):

@satellite73

OpenStudy (johnweldon1993):

 Looks a lot like a right triangle doesn't it :)

OpenStudy (anonymous):

yes it is how to solve this easily?

OpenStudy (johnweldon1993):

Well the Pythagorean theorem comes into play since this is a right triangle. We have 2 legs of the triangle and want to know the hypotenuse.. \[\large a^2 + b^2 = c^2\] We know 'c' is the hypotenuse so we do \[\large c = \sqrt{a^2 + b^2}\] Just plug in 100 for 'a' and 500 for 'b' and solve for 'c'

OpenStudy (anonymous):

so i will only substitute their respective values

OpenStudy (anonymous):

answer is 509.9

OpenStudy (johnweldon1993):

That is correct, the resultant would be 509.9 m/s in the northeast direction

OpenStudy (anonymous):

okay thanks a lot